OpenStudy (anonymous):

Solve for u u^2-3u-4=0

OpenStudy (jahnavi):

\[u ^{2} -3u -4 = 0\] \[u ^{2} -4u + 1u - 4 = 0\] \[u (u - 4)+ 1 ( u - 4) = 0\] \[(u + 1) (u - 4) = 0\] therefore, \[u = -1 u = +4\]